# Strategic Classification with Non-Linear Classifiers

### v1

@@ 4 Points
For linear $h$, the point mapping is always one-to-one. For non-linear $h$, the mapping remains one-to-one as long as $h$ is ‘well-behaved’ on $x$. We consider this the ‘typical’ case and note that each $x$ is mapped to the single point ${\hat{x}}=x-\alpha\hat{n}_{x}$, where $\hat{n}_{x}$ is the unit normal at $x$ w.r.t. $h$ (see Appendix B.1). A necessary condition for one-to-one mapping is that $h$ is smooth at $x$, but smoothness is not sufficient—as the following cases depict.
@@
Even if $h$ is smooth at $x$, in some cases the point mapping will be null, $x\mapsto\varnothing$. We refer to this phenomenon as wipeout, and identify two types. Direct wipeout occurs when the signed curvature $\kappa$ at $x$ is negative and of sufficiently large magnitude; this is defined formally in Sec. 5.1. Intuitively, this results from strategic behavior obfuscating any rapid changes in the decision boundary of $h$. Indirect wipeout occurs when the the set of points $\nabla_{h}(x)$ is contained in $B_{\alpha}(x^{\prime})$ for some other $x^{\prime}$ on the decision boundary. For both types, the result is that $B_{\alpha}(x)$ is entirely contained in the positive region of $h$ and such $x$ have no effect on $h_{\Delta}$ (see Appendix B.2).
@@
When the decision boundary of $h$ at $x$ is not smooth (i.e., a kink or corner), $x$ can be mapped to a continuum of points on the effective boundary of $h_{\Delta}$. This, however, occurs only when the decision boundary is locally convex towards the positive region. The set of mapped points $\nabla_{h}(x)$ are those on $S_{\alpha}(x)$ who do not intersect with any other $B_{\alpha}(x^{\prime})$ for other $x^{\prime}$ on the boundary of $h$. Thus, the shape of the induced ‘artifact’ is partial to a ball. Partial expansion may occur if some of the expansion artifact is discarded through indirect wipeout (Appendix C.1).
@@
A final case is when a set of points $\{x\}$ on the boundary of $h$ are all mapped to the same single point ${\hat{x}}$ on $h_{\Delta}$, i.e., $\nabla_{h}(x)=\nabla_{h}(x^{\prime})={\hat{x}}$ for all $x,x^{\prime}$ in the set. Typically, this happens at the transitions between well-behaved decision boundary segments and wiped-out regions. Collision often results in non-smooth kinks on $h_{\Delta}$, even if all source points were smooth on $h$.
5 Classifiers
We now turn to examine how strategic behavior affects the classifier mapping $h\mapsto h_{\Delta}$. We identify two mechanisms that can limit expressivity of effective classifiers: curvature, and containment.
@@ 5.1 Curvature
To understand the mapping $h\mapsto h_{\Delta}$, it will be useful to first examine what happens to boundary points $x$ in terms of how $h$ behaves around them. We make use of the notion of signed curvature, denoted $\kappa$, which quantifies how much the decision boundary deviates from being linear at a certain point $x$. For $d=2$, curvature is formally defined as $\kappa=1/r$ where $r$ is the radius of the circle that best approximates the curve at $x$. For $d>2$, we use $\kappa^{*}=\sup(K)$ where $K$ includes all directional curvatures at $x$ (see Appendix A.1). The sign of $\kappa$ measures whether the decision boundary bends (i.e., is convex) towards the positive region ($\kappa>0$), the negative region ($\kappa<0$), or is locally linear ($\kappa=0$).
Let $x$ be a typical point on the decision boundary of $h$ which satisfies the one-to-one mapping detailed in Sec. 4. Our next result (which relies on an equivalent form in the field of offset curves [12]) characterizes the effective signed curvature of $h_{\Delta}$ at ${\hat{x}}$, denoted $\kappa_{\Delta}$, as a function of the curvature $\kappa$ of $h$ at $x$.
@@
Let $x$ be a point on the decision boundary of $h$ with signed curvature $\kappa\geq-1/\alpha$. The effective curvature of the corresponding ${\hat{x}}$ on the boundary of $h_{\Delta}$ is given by:
@@
Figure 3: (Top) Graph of Eq. 3 with asymptotes represented as dashed lines. (Bottom) Depictions of curvature mappings for increaseing $\kappa$.
When $d>2$, each curvature in $K$ is mapped as in Eq. 3 to form $K_{\Delta}=\{\kappa/(1+\alpha\kappa):\kappa\in K\}$, and $\kappa^{*}_{\Delta}$ follows. See Appendix B.3 for proof and Fig. 3 for an illustration of how this impacts the shape of $h_{\Delta}$. Prop. 1 entails several interesting properties. First, strategic behavior preserves the direction of curvature, i.e., $\kappa\cdot\kappa_{\Delta}\geq 0$ always. In particular, $\kappa=0$ entails $\kappa_{\Delta}=0$, which explains why linear functions remain linear. Second, effective curvature $\kappa_{\Delta}$ is monotonically increasing in $\kappa$, concave, and sub-linear. This implies that for $\kappa>0$ the effective positive curvature decreases, while for $\kappa<0$ the effective negative curvature increases. Thus, strategic behavior acts as a directional smoothing operator, i.e., the boundaries’ rate of change is expected to decrease but only in the positive direction.
@@
When $d>2$, Obs. 1 holds for each curvature $\kappa\in K$ and can be summarized as $\kappa^{*}_{\Delta}\leq 1/\alpha$. Though Prop. 1 only implies Obs. 1 at points on $h_{\Delta}$ that are typical, we will show in Sec. 5.3 that is also true at any point on $h_{\Delta}$. The above holds irrespective of the shape and curvature of the original $h$. In particular, non-smooth points on $h$ (i.e., with $\kappa\to\infty$), such as an edge or vertex at the intersection of halfspaces, becomes smooth on $h_{\Delta}$. Conversly, in the negative direction, $\kappa_{\Delta}$ approaches $\infty$- ∞ as $\kappa\to 1/\alpha$. Thus, negative curvature can grow arbitrarily large, and at $\kappa=-1/\alpha$ maps to a non-smooth "kink" in $h_{\Delta}$. $\kappa_{\Delta}$ is no longer defined when $\kappa$ crosses the asymptote at $-1/\alpha$ because such points are wiped out and have no corresponding ${\hat{x}}$ on the boundary of $h_{\Delta}$. This is the exact case of indirect wipeout in Sec. 4 (see Appendix B.2 for proof).
@@
Any point $x$ on the decision boundary of $h$ with curvature $\kappa<-1/\alpha$ has no influence on $h_{\Delta}$. In higher dimensions, this is true for any point $x$ with $\inf(K)<-1/\alpha$.
@@ 5.2 Containment
Because the decision boundary is a level set of some score function $f$, contained regions are not only possible, but realistic. Consider a given classifier $h$ where the decision boundary forms some bounded connected negative region $C\subset\mathbb{R}^{d}$ where $h(x)=0\,\,\forall x\in C$. Strategic behavior dictates that negative areas shrink, and as such, only a subset of $C$ will still be classified as negative in $h_{\Delta}$. If $C$ is sufficiently small, then it will disappear completely leaving $h_{\Delta}(x)=1\,\,\forall x\in C$.
@@
Any negative region $C$ fully contained in some $B_{\alpha}(x)$ will become positive in $h_{\Delta}$.
The fact that positive areas grow and negative areas shrink also implies that there is a minimum size positive region that can exist in any $h_{\Delta}$. This idea will be formalized and generalized later in Sec. 5.3. Another observation is that when negative regions are wiped out, two or more positive regions may merge leading to a less expressive decision boundary. On the other hand, the strategic setting can also split a negative region creating a more expressive decision boundary. For example, if $h$ induces a negative connected region $C$ whose shape is formed by two large regions connected by a narrow "strait", the strait will be wiped out leaving two disjoint components (see Appendix C.2).
@@
Figure 4: Types of impossible effective classifiers. From left to right: small positive region, narrow positive strip, large positive curvature, piecewise (hyper)linear convex towards positive region.
5.3 Impossible effective boundaries
Based on the above, we derive two general statements about effective classifiers that cannot exist. Each statement considers some candidate classifier $g$, and shows that if $g$ satisfies a certain property, then there is no classifier $h$ for which $g$ is its induced effective classifier, i.e., $g\neq h_{\Delta}$ for any $h$.
We require some additional notation. Given $h$, define the set of negatively classified points as ${\cal{X}}_{0}(h)=\{x\,:\,h(x)=0\}$. For a set $A\subseteq{\cal{X}}$, we say a point $x$ is reachable from $A$ if $\exists x^{\prime}\in A\,\,:c(x,x^{\prime})\leq\alpha$.
@@ 5.3 Impossible effective boundaries
Let $g$ be a candidate classifier. If there exists a point $x$ on the decision boundary of $g$ such that all points $x^{\prime}\in S_{\alpha}(x)$ are reachable from some point in ${\cal{X}}_{0}(g)$, then there is no $h$ such that $g=h_{\Delta}$.
The intuition behind Prop. 2 is that $g$ cannot simultaneously classify $x$ as positive while also classifying all $y\in{\cal{X}}_{0}$ as negative. Full proof in Appendix B.4. As stated in Prop. 3, when $g$ is smooth at $x$, Prop. 2 can be simplified to require that only a single point be checked instead of all of $S_{\alpha}(x)$.
@@
Let $g$ be a candidate classifier. If there exists a point $x$ on the decision boundary of $g$ where (i) $g$ is smooth, and (ii) the point ${\hat{x}}=x+\alpha\hat{n}_{x}$ is reachable by some other $x^{\prime}\in{\cal{X}}_{0}(g)$, then there is no $h$ such that $g=h_{\Delta}$.
The intuition behind Prop. 3 is the same as for Prop. 2 except that when $g$ is smooth at $x$, the only point in $S_{\alpha}(x)$ that may not be reachable by some $x^{\prime}\in{\cal{X}}_{0}(g)$ is ${\hat{x}}=x+\alpha\hat{n}_{x}$ (see Appendix B.5).
@@
The implication of Props. 2 and 3 is that there exist decision boundaries which are generally realizable, but are no longer realizable as effective classifiers in the strategic setting. Note that because it only takes a single decision boundary point to render an entire effective classifier infeasible, each of these examples are broad categories and pose meaningful restrictions on the set of possible effective classifiers. We highlight four types of impossible effective decision boundaries:
@@
These are illustrated in Fig. 4. Proofs in Appendix C.3, and either build on Obs. 1-4, or show the conditions of Props. 2 or 3 hold. From these four example, we note the following observation.
@@
In Appendix C.4 we show that there are actually infinite $h$ that map to nearly all possible $h_{\Delta}$. Furthermore, Obs. 5 will be important to proving Thm. 1 and Cor. 1
6 Model classes
On the classifier level, the strategic setting is clearly more restrictive than the regular setting. Indeed, in Sec. 6.2 we will extend this idea into the model class level by analyzing the impacts on universality and optimal accuracy. However, as we show in 6.1 the model class mapping is more nuanced than the classifier mapping since complexity can actually increase as a result of strategic behavior.
@@ 6.1 Class complexity
Recall that for linear classes, $H_{\Delta}$ remains linear and so $H\mapsto H_{\Delta}=H$. Although this equivalence is not unique to the class of linear classifiers (see examples in Appendix C.5), typically, we can expect strategic behavior to alter the complexity of the induced class. Here, we use VC analysis to shed light on this effect by examining the possible relations between $\mathrm{VC}(H)$ and $\mathrm{VC}(H_{\Delta})$. [^1] This aligns with the notion of strategic VC used in [31, 21], i.e., $\mathrm{SVC}(H)\equiv\mathrm{VC}(H_{\Delta})$.
@@
Since our results so far point out that effective classifiers are inherently constrained, a plausible guess is that effective complexity should decrease. Indeed, there are some natural classes where VC cannot increase. For example:
@@
For each class $Q^{k}_{s}$ of negative-inside $k$-vertex polytopes bounded by a radius of $s$, $\mathrm{VC}(Q^{k}_{s})\geq\mathrm{VC}(Q_{s,\Delta}^{k})$. Moreover, if $s=\infty$, then $\mathrm{VC}(Q^{k}_{s})=\mathrm{VC}(Q_{s,\Delta}^{k})$ as implied by Thm. 2.
Proof in Appendix B.7. Curiously, in some extreme cases, the reduction spans the maximum extent.
@@
Proof in Appendix B.8. See Appendix C.6 for other examples where VC strictly decreases yet does not diminish completely. Thm. 1 carries concrete implications for strategic learning:
@@
A model class that is not learnable in the regular setting may become learnable in strategic environments.
@@
While lower effective VC is indeed a possibility, the effective complexity generally stays the same or increases. As an intuition for how the VC might increase, consider the simple example in Fig. 5 (left). Here $H$ includes 4 non-overlapping classifiers with $\mathrm{VC}(H)=1$, yet under strategic behavior, the effective classifiers overlap increasing the shattering capacity to 2. This case can be generalized to give $VC(H_{\Delta})=\theta(d\cdot VC(H))$ (see Appendix C.7). We next identify a simple generic sufficient condition for the effective complexity of any learnable class to be non-decreasing. The proof of Thm. 2 can be found in Appendix B.9.
@@
This applies to many common classes, such as polynomials, polytopes, and most neural networks. In the case of polytopes (positive-inside), we can also upper bound the VC of the effective class.
@@
Consider either the $\ell_{1}$, $\ell_{2}$, or $\ell_{\infty}$ costs and let $H\subseteq P^{k}$, where $P^{k}$ is the set of all $k$-vertex polytopes in $R^{d}$. Then $\mathrm{VC}(H_{\Delta})=O(kd^{2}\log k)$.
Proof is in Appendix. B.10. Since it is known that $\mathrm{VC}(P^{k})=O(kd^{2}\log k)$ [22], Thm. 3 suggests that the effective VC dimension can increase, but potentially only to a limited extend as in this case where it remains in the same order. For $\ell_{1}$ and $\ell_{\infty}$, note the effective class remains within the polytope family, but with additional vertices (see Appendix B.6). For $\ell_{2}$, the bound holds despite $H_{\Delta}$ no longer including only polytopes (since corners in $h$ can become "rounded" in $h_{\Delta}$).
As a follow-up to Thm. 3 and Cor. 1, it is natural to wonder if there is a maximum to how much $VC(H_{\Delta})$ can grow? Despite our best efforts, we could not find any class with bounded VC whose effective VC becomes unbounded, and therefore suspect such classes do not exist.
Conjecture 1.
It is likely that for any norm cost, if $H$ is learnable, then $H_{\Delta}$ is also learnable.
If proven, this would extend the result of Cohen et al. [6] in the special case of finite manipulation graphs to the more common setting of explicit cost functions (where the graph is continuous). [^2] In particular, they show that $\mathrm{VC}(H_{\Delta})=\tilde{\Theta}(\mathrm{VC}(H)\cdot\log k)$ where $k$ is the degree of the manipulation graph. Note the bound is vacuous for $k\to\infty$, and the proof technique does not apply to continuous graphs.

### v2

@@ 4.1 Point mapping
For linear $h$, the point mapping is always one-to-one. For non-linear $h$, the mapping remains one-to-one as long as $h$ is ‘well-behaved’ around z{z}. We consider this the ‘typical’ case and note that each z{z} is mapped to its $\alpha$-offset point ${x}={z}-\alpha\hat{n}_{{z}}$, where $\hat{n}_{{z}}$ is the unit normal at z{z} w.r.t. $h$ (see Appendix B.1). A necessary condition for one-to-one mapping is that $h$ is smooth at z{z}, but smoothness is not sufficient, as the following cases depict.
@@
Even if $h$ is smooth at z{z}, in some cases the point mapping will be null: ${z}\mapsto\varnothing$. We refer to this phenomenon as wipeout, and identify two types. Direct wipeout occurs when the signed curvature $\kappa$ at z{z} is negative and of sufficiently large magnitude; this is defined formally in Sec. 4.2. Intuitively, this results from strategic behavior obfuscating any rapid changes in the decision boundary of $h$. Indirect wipeout occurs when the set of points $\nabla_{h}({z})$ is contained in $B_{\alpha}(x^{\prime})$ for some other $x^{\prime}$ on the decision boundary. For both types, the result is that $B_{\alpha}({z})$ is entirely contained in the positive region of $h_{\Delta}$ and such z{z} have no effect on $h_{\Delta}$ (see Appendix B.2).
@@
When the decision boundary of $h$ at z{z} is not smooth (i.e., a kink or corner), z{z} can be mapped to a continuum of points on the effective boundary of $h_{\Delta}$. This, however, occurs only when the decision boundary is locally convex towards the positive region. The set of mapped points $\nabla_{h}({z})$ are those on $S_{\alpha}({z})$ which do not intersect with any other $B_{\alpha}(x^{\prime})$ for any other $x^{\prime}$ on the boundary of $h$. Thus, the shape of the induced ‘artifact’ is partial to a ball. Partial expansion may occur if some of the expansion artifact is discarded through indirect wipeout (Appendix C.2).
@@
A final case is when a set of points $\{{z}_{1}...{z}_{n}\}$ on the boundary of $h$ are all mapped to the same single point x{x} on $h_{\Delta}$, i.e., $\nabla_{h}({z})=\nabla_{h}({z}^{\prime})={x}$ for all ${z},{z}^{\prime}$ in the set. Typically, this happens at the transitions between well-behaved decision boundary segments and wiped-out regions. Collision often results in non-smooth kinks on $h_{\Delta}$, even if all source points were smooth on $h$.
@@ 4.2 Curvature
Building up to the level of classifiers, it is first useful to examine what happens to boundary points z{z} in terms of how $h$ behaves around them. We make use of the notion of signed curvature, denoted $\kappa$, which quantifies how much the decision boundary deviates from being linear at a certain point $x$. For $d=2$, curvature is formally defined as $\kappa=1/r$, where $r$ is the radius of the circle that best approximates the curve at $x$. For $d>2$, we use $\kappa^{*}=\sup(K)$, where $K$ includes all directional curvatures at $x$ (see Appendix A.1). The sign of $\kappa$ measures whether the decision boundary bends (i.e., is convex) towards the positive region ($\kappa>0$), the negative region ($\kappa<0$), or is locally linear ($\kappa=0$).
Let z{z} be a typical point on the decision boundary of $h$ which satisfies the one-to-one mapping detailed in Sec. 4.1. We borrow a classic result from the field of offset curves [12] to calculate the effective signed curvature of $h_{\Delta}$ at x{x}, denoted $\kappa_{\Delta}$, as a function of the curvature $\kappa$ of $h$ at z{z}.
@@
Let z{z} be a point on the decision boundary of $h$ with signed curvature $\kappa\geq-1/\alpha$. The effective curvature of the corresponding x{x} on the boundary of $h_{\Delta}$ is given by:
@@
When $d>2$, each $\kappa\in K$ is mapped as in Eq. 3 to form $K_{\Delta}=\{\kappa/(1+\alpha\kappa):\kappa\in K\}$, and $\kappa^{*}_{\Delta}$ follows. See Appendix B.3 for proof and an illustration of how this impacts the shape of $h_{\Delta}$. Prop. 1 entails several interesting properties. First, strategic behavior preserves the direction of curvature, i.e., $\kappa\cdot\kappa_{\Delta}\geq 0$ always. In particular, $\kappa=0$ entails $\kappa_{\Delta}=0$, which explains why linear functions remain linear. Second, the effective curvature $\kappa_{\Delta}$ is monotonically increasing in $\kappa$, concave, and sub-linear. This implies that for $\kappa>0$ the effective positive curvature decreases, while for $\kappa<0$ the effective negative curvature increases. Thus, strategic behavior acts as a directional smoothing operator, i.e., the boundaries’ rate of change is expected to decrease but only in the positive direction.
@@
When $d>2$, Obs. 1 holds for each curvature $\kappa\in K$ and can be summarized as $\kappa^{*}_{\Delta}\leq 1/\alpha$. Though Prop. 1 only implies Obs. 1 at points on $h_{\Delta}$ that are typical, we will show in Sec. 4.4 that it is also true at any point on $h_{\Delta}$. The above holds irrespective of the shape and curvature of the original $h$. In particular, non-smooth points on $h$ (i.e., with $\kappa\to\infty$), such as an edge or vertex at the intersection of halfspaces, become smooth on $h_{\Delta}$. Conversely, in the negative direction, $\kappa_{\Delta}$ approaches -$\infty$ as $\kappa\to 1/\alpha$. Thus, negative curvature can grow arbitrarily large, and at $\kappa=-1/\alpha$ maps to a non-smooth "kink" in $h_{\Delta}$. $\kappa_{\Delta}$ is no longer defined when $\kappa$ crosses the asymptote at $1/\alpha$ because such points are wiped out and have no corresponding x{x} on the boundary of $h_{\Delta}$. This is the exact case of indirect wipeout in Sec. 4.1 (see Appendix B.2 for proof and Appendix C.10 for practical use in learning).
@@
Any point z{z} on the decision boundary of $h$ with curvature $\kappa<-1/\alpha$ has no influence on $h_{\Delta}$. In higher dimensions, this is true for any point $x$ with $\inf(K)<-1/\alpha$.
@@ 4.3 Containment
Because the decision boundary is a level set of some (non-linear) score function $f$, contained regions are not only possible, but realistic. Consider a given classifier $h$ where the decision boundary forms some bounded connected negative region $C\subset\mathbb{R}^{d}$ where $h(x)=0\,\,\forall x\in C$. Strategic behavior dictates that negative areas shrink, and so only a subset of $C$ will still be classified as negative in $h_{\Delta}$. If $C$ is sufficiently small, then it will disappear completely leaving $h_{\Delta}(x)=1\,\,\forall x\in C$.
@@
Any negative region $C$ fully contained in some $B_{\alpha}({x^{\prime}})$ will become positive in $h_{\Delta}$.
The fact that positive areas grow and negative areas shrink also implies that there is a minimum positive region size that can exist in any $h_{\Delta}$. This idea will be formalized and generalized later in Sec. 4.4. One implication of Obs. 3 is that when negative regions are wiped out, two or more positive regions may merge, leading to a less expressive decision boundary. On the other hand, the strategic setting can also split a negative region, creating a more expressive decision boundary (see Appendix C.3).
@@
Figure 3: Types of impossible effective classifiers. From left to right: small positive region, narrow positive strip, large positive curvature, piecewise (hyper)linear convex towards the positive region.
4.4 Impossible Effective Classifiers
Based on the above, we derive two general statements about effective classifiers that cannot exist. Each statement considers some candidate classifier $g$, and shows that if $g$ satisfies a certain property, then there is no classifier $h$ for which $g$ is its induced effective classifier, i.e., $g\neq h_{\Delta}$ for any $h$. We first require some additional notation. Given $h$, define the set of negatively classified points as ${\cal{X}}_{0}(h)=\{x\,:\,h(x)=0\}$. For a set $A\subseteq{\cal{X}}$, we say a point $x$ is reachable from $A$ if $\exists x^{\prime}\in A\,\,:c(x,x^{\prime})\leq\alpha$.
@@ 4.4 Impossible Effective Classifiers
Let $g$ be a candidate classifier. If there exists a point x{x} on the decision boundary of $g$ such that all points $x^{\prime}\in S_{\alpha}({x})$ are reachable from some point in ${\cal{X}}_{0}(g)$, then there is no $h$ such that $g=h_{\Delta}$.
The intuition behind Prop. 2 is that no $h$ can simultaneously strategically classify x{x} as positive while all $y\in{\cal{X}}_{0}(g)$ as negative. Full proof in Appendix B.4. As stated in Prop. 3, when $g$ is smooth at x{x}, Prop. 2 can be simplified to require that only a single point be checked instead of all of $S_{\alpha}({x})$.
@@
Let $g$ be a candidate classifier. If there exists a point x{x} on the decision boundary of $g$ where (i) $g$ is smooth, and (ii) the offset point ${\hat{x}}={x}+\alpha\hat{n}_{x}$ is reachable by some other $x^{\prime}\in{\cal{X}}_{0}(g)$, then there is no $h$ such that $g=h_{\Delta}$.
The intuition behind Prop. 3 is the same as for Prop. 2, except that when $g$ is smooth at x{x}, the only point in $S_{\alpha}({x})$ that may not be reachable by some $x^{\prime}\in{\cal{X}}_{0}(g)$ is the $\alpha$-offset (see Appendix B.5).
@@
The implication of Props. 2 and 3 is that there exist decision boundaries that are generally realizable, but are no longer realizable as effective classifiers in the strategic setting. Note that because it only takes a single decision boundary point to render an entire effective classifier infeasible, each of these examples are broad categories and pose meaningful restrictions on the set of possible effective classifiers. As illustrated in Fig. 3, we highlight four types of impossible effective decision boundaries:
@@
All proofs are in Appendix C.4, and either build on Obs. 1-4, or show the conditions of Props. 2 or 3 hold. From these four examples, we note the following observation.
@@
In Appendix C.5 we show that there are actually infinite $h$ that map to nearly all possible $h_{\Delta}$. Furthermore, Obs. 5 will be important for proving Thm. 1 and Cor. 1.
5 Class-level analysis
On the classifier level, the strategic setting is clearly more restrictive than the regular setting. Indeed, in Sec. 5.2 we extend this idea into the model class level by analyzing the impacts on universality and optimal accuracy. However, as we show in Sec. 5.1, the model class mapping is more nuanced than the classifier mapping since complexity can actually increase as a result of strategic behavior.
@@ 5.1 Class complexity
Recall that for linear classes, $H_{\Delta}$ remains linear, and so $H\mapsto H_{\Delta}=H$. Although this equivalence is not unique to the class of linear classifiers (see examples in Appendix C.6), typically, we can expect strategic behavior to alter the complexity of the induced class. Here, we use VC analysis to shed light on this effect by examining the possible relations between $\mathrm{VC}(H)$ and $\mathrm{VC}(H_{\Delta})$. [^1] This aligns with the notion of strategic VC used in [33, 22], i.e., $\mathrm{SVC}(H)\equiv\mathrm{VC}(H_{\Delta})$.
@@
Since our results so far point out that effective classifiers are inherently constrained, a plausible guess is that effective complexity should decrease. Indeed, there are some natural classes where the VC cannot increase. For example:
@@
Let $Q^{k}_{s}$ be the class of negative-inside $k$-vertex polytopes bounded by a radius of $s$, then $\mathrm{VC}(Q^{k}_{s})\geq\mathrm{VC}(Q_{s,\Delta}^{k})$. Moreover, if $s=\infty$, then $\mathrm{VC}(Q^{k}_{s})=\mathrm{VC}(Q_{s,\Delta}^{k})$ as implied by Thm. 2.
The proof for Prop. 4 can be found in Appendix B.7. Curiously, in some extreme cases, the VC reduction spans the maximum extent.
@@
The proof for Thm. 1 can be found in Appendix B.8. See Appendix C.7 for other examples where VC strictly decreases but does not diminish completely. Thm. 1 carries concrete implications for strategic learning:
@@
Some model classes that are not learnable become learnable in strategic environments.
@@
Though a lower effective VC is indeed a possibility, a more plausible outcome is for it to stay the same or increase. As an intuition for how the VC might increase, consider the simple example in Fig. 4 (left). Here, $H$ includes four non-overlapping classifiers (each color is a separate classifier) with $\mathrm{VC}(H)=1$. Yet, with strategic behavior, the effective classifiers (dashed curves) overlap, increasing the shattering capacity to 2. This case can be generalized to give $VC(H_{\Delta})=\theta(d\cdot VC(H))$ (see Appendix C.8). We next identify a simple generic sufficient condition for the effective complexity of any learnable class to be non-decreasing.
@@
The proof for Thm. 2 can be found in Appendix. B.9. Thm. 2 applies to many common classes, including polynomials, piecewise linear functions, and most neural networks. See Appx. C.10 for implications on model selection.
Because Thm. 2 suggests that the effective VC can increase, but does not state by how much, one concern would be that this increase can be unbounded, which would deem the effective class unlearnable. Our next result shows that this is not the case even for the common and highly expressive class of piecewise-linear classifiers (with a bounded number of segments), which includes neural networks with ReLU activations (of fixed size) as a special case, and is often used in studies on neural network approximation [e.g., 28, 19].
@@
Let $H^{m,k}$ be a class of piecewise-linear classifiers with at most m segments and k intersections of linear segments. Then $\mathrm{VC}(H_{\Delta}^{m,k})=O\big(d\,m\,log(m)+\nu_{p}\,k\,log(k)\big)$, where $\nu_{p}$ is the VC of the class of $\ell_{p}$ norm balls, and is $\Theta(d)$ for $p=1,2,\infty$.
The proof for Thm. 3 is in Appendix B.10. We also prove that $\mathrm{VC}(H_{m,k})=O\big(d\,m\,log(m)\big)$, suggesting that the effective VC dimension can increase, but only to a limited extent. The main implication of Thm. 3 is therefore that under these conditions, strategic behavior maintains learnability: for any $H$ that can be expressed as (or approximated by) piecewise-linear functions, if $H$ (or its approximation) is learnable, then $H_{\Delta}$ is also learnable. We conjecture that this relation remains true for general $H$; see Appendix A.3 for discussion on the challenges of the general case and connections to [6]. For the special case of positive-inside polytopes — a subset of the piecewise-linear class — the effective VC dimension even maintains the same order of magnitude.
Theorem 4.
[…]
The proof for Thm. 4 is in Appendix. B.11. Note that for the $\ell_{1}$ and $\ell_{\infty}$ costs, the effective class remains within the polytope family, but potentially with additional vertices (see Appendix B.6). For the $\ell_{2}$ cost, the bound still holds despite $H_{\Delta}$ no longer including only polytopes (since corners in $h$ can become "rounded" in $h_{\Delta}$).